Python))

  1. A copy center charges 5 cents per copy for the first 100 copies and 3 cents per copy for each additional copy. Write a program that requests the number of copies as input and display the total cost.
  2. Federal law requires that hourly employees be paid “Time and a half” for work in excess of 40 hours in a week. For example, if a person’s hourly wage is $ 12 and he/ she works 60 hours in a week, the person’s gross pay should be:

             (40 * 12) + (1.5 * 12* (60 – 40)) = $840
Answer #1

Answer 1:

copies = float(input("Enter number of copies : "))

total=0

if(copies<100):

total=copies*5

else:

total=100*5

total=total+(copies-100)*3

total=total/100

print("Total : $",total)

Answer 2:

h = float(input("Enter number of hours : "))

total=0

if(h<40):

total=5*12

else:

total=40*12

total=total+(h-40)*18


print("Total : $",total)